OpenStudy (omgthatgirl):

What is the solution of the equation X^2+23=0 A. X=-6+sqrt59 B. X=-6+sqrt23 C. X=-6+sqrt13 D. X=-6+i sqrt13

OpenStudy (welshfella):

are you sure you have the correct equation?

OpenStudy (omgthatgirl):

Oh sorry it's x^2+12x+23=0

OpenStudy (omgthatgirl):

@welshfella

OpenStudy (welshfella):

Ok - do you know the formula for solving quadratic equations?

OpenStudy (omgthatgirl):

Something like ax+b...?

OpenStudy (welshfella):

for the equation ax^2 + bx + c = 0 the formula is x = [ -b +/- sqrt(b^2 - 4ac)] / 2a

OpenStudy (omgthatgirl):

K..

OpenStudy (welshfella):

if you compare this equation with that in the equation:- ax^2 + bx + c = 0 and x^2 +12x + 23 = 0 you can see which values of a , b and c you require so a = 1, b = 12 and c = ? cam you see what c is?

OpenStudy (welshfella):

* can

OpenStudy (omgthatgirl):

23

OpenStudy (welshfella):

right so insert them in the formula i'll start you off

OpenStudy (welshfella):